Solve
a) 7p^(-1) + 2p^(-2) - 4 = 0
Can you please help me out? Thanks so much in advance:)

Answer by lwsshak3(11628)

Solve
a) 7p^(-1) + 2p^(-2) - 4 = 0
let p^(-1)=x
p^(-2)=x^2
..
2x^2+7x-4=0
(2x-1)(x+4)=0
..
x=1/2=p^(-1)=1/p
p=2
..
x=-4=p^(-1)=1/p
p=-1/4